OpenStudy (anonymous):

Evaluate each function for the given value. f(g-1) if f(x) = x2 -2x +1

OpenStudy (matheducatormcg):

If that is \[f(g-1)\] then \[f(g-1)=(g-1)^2-2(g-1)+1\] which expands to \[f(g-1)=g^2-2g+1-2g+2+1\] then combine like terms \[f(g-1)=g^2-4g+4\]